OpenStudy (pokemon23):

solve for x in terms of y: 2x-y=8 show all work solve for x in terms of y: 2x-y=8 show all work @Mathematics

OpenStudy (saifoo.khan):

x = (y + 8)/2

OpenStudy (anonymous):

isolate the xs.

OpenStudy (pokemon23):

guys I'm retarded speak in english

OpenStudy (pokemon23):

show me the steps

OpenStudy (agreene):

2x-y=8 add y to both sides 2x=8+y divide by 2 x=(8+y)/2 separate fractions x=8/2+y/2 evaluate fractions \[x=2+\frac{y}{2}\]